OpenStudy (anonymous):

Write the equation y=-x^2-3x in vertex form. what is the vertex and the axis of symmetry

OpenStudy (cwrw238):

vertex form = (x - 1.5)^2 - 1.5^2

OpenStudy (anonymous):

how did ou get it ? and do you think you can help me with some more ? please

OpenStudy (cwrw238):

to convert a quadratic to vertex form you use the identity x^2 + bx + c = (x + b/2)^2 - (b/c)^2 + c in your case b was = -3 so b/2 = -1.5 and c was = 0